The sum of two numbers is 25. One number is 7 less than 3 times the other number.
x=larger number
y=lesser number
This is what I have so far. What do I do next?
x=9
x+y=120

Answer by checkley77(12844)

x=9
x+y=120
QUESTION?
WHERE DID YOU GET X=9 & X+Y=120???????
-------------------------------------
X+Y=25 OR X=25-Y
X=3Y+7
25-Y=3Y+7
-Y-3Y=7-25
-4Y=-18
Y=-18/-4
Y=4.5 ANS. FOR THE SMALLER NUMBER.
X=25-4.5
X=20.5 ANS. FOR THE LARGER NUMBER.
PROOF:
20.5=3*4.5+7
20.5=13.5+7
20.5=20.5